OpenStudy (anonymous):

Let k be an integer greater or equal to zero, show that \[ \int_0^{\pi } \cos ^{2 k+1}(x) \, dx=0 \\ \int_0^{\pi } \cos ^{2 k}(x) \, dx=\frac{\pi (2 k)!}{2^{2 k} (k!)^2} \]

OpenStudy (anonymous):

Hint, read http://en.wikipedia.org/wiki/Integration_by_reduction_formulae

OpenStudy (anonymous):

\[Let f(x)=\left( \cos x \right)^{2k+1}\] \[f \left( \pi-x \right)=\left\{ \cos \left( \pi-x \right) \right\}^{2k-1}=\left( -\cos x \right)^{2k+1}=-\left( \cos x \right) ^{2k+1}=-f \left( x \right)\] \[\int\limits_{0}^{\pi}f(x)=0\]

ganeshie8 (ganeshie8):

\[\begin{align}\int_0^{\pi } \cos ^{2 k}(x) \, dx&= \frac{2k-1}{2k} \int_0^{\pi} \cos^{2k-2}(x) \,dx\\~\\ &=\frac{2k-1}{2k} \cdot \frac{2k-3}{2k-2}\cdots \int_0^{\pi} 1\, dx \end{align}\]

OpenStudy (anonymous):

This result was used to prove the link above. @SithsAndGiggles